OpenStudy (anonymous):

Choose the correct simplification of 3^8/3^3

OpenStudy (anonymous):

243 15 1/43 1/15

OpenStudy (ajprincess):

\[\Large\frac{ a^b }{ a^c }=a^{b-c}\]

OpenStudy (anonymous):

i dnt get it

OpenStudy (mathstudent55):

When you divide a power by a power, and they have the same base, write the same base and subtract the exponents. That's the rule @ajprincess wrote above. Example: 7^5/7^3 You are dividing two powers, 7^5 and 7^3. They have the same base, 7, so the answer is: write the same base, 7, and the power is the subtraction of the exponents: 5 - 3 = 2. Answer is 7^2 = 7*7 = 49

OpenStudy (anonymous):

so the answer is either 15 or 243 @mathstudent55

OpenStudy (anonymous):

i got 243 though

OpenStudy (imtant):

3^8-^3=3^5 So it is 243.

OpenStudy (mathstudent55):

3^8/3^3 = 3^(8-3) = 3^5 = 3*3*3*3*3 = 243